Circular No. 6228

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

COMET 1995 S1 Word has been received from Japan of the independent visual discov-

eries by Yuji Nakamura, Masaaki Tanaka, and Shougo Utsunomiya of a comet, with the following information available:

1995 UT  α_{2000}  δ_{2000}  m_{1} Observer
Sept.17.792  9^{h}03.^{m}5 − 2^{o}28′ 7.0 Nakamura
17.802 9 03 − 2 30 7 Utsunomiya
17.80635 9 03.4 − 2 27  Tanaka

Y. Nakamura (Suzuka, Mie). 20×120 binoculars. Comet diffuse; coma diameter 2′. Motion noted as 1′.5 northward in 50 min. Communicated by T. Nakamura, National Astronomical Observatory, Tokyo.

S. Utsunomiya (Minamioguni, Kumamoto). 25×150 binoculars. Coma diameter 5′; central condensation present. Possible slight eastward motion in 40 min. Communicated by A. Nakamura, Kuma, Ehime.

M. Tanaka (Iwaki, Fukushima). Discovered on Sept. 17.795 with 25×150 binoculars, visual m_{1} = 7.5, coma diameter 5′, strong central condensation. Above position was from a 2-min exposure with a 0.20-m f/1.5 Schmidt camera (+ TP2415 film) that shows an ion tail 25′ long in p.a. 260^{o} and a dust tail 5′ long in p.a. 320^{o}. Communicated by T. Hirayama, National Astronomical Observatory, Tokyo.

COMET 6P/D’ARREST M. J. Mumma, Goddard Space Flight Center (GSFC), NASA; M. A.

DiSanti, Universities Space Research Association and GSFC; and X. Xie, National Academy of Sciences, National Research Council, and GSFC, report the detection of water vapor in 6P/d’Arrest using the NASA Infrared Telescope Facility (+ CSHELL) with the comet at r = 1.42 AU and Δ = 0.47 AU: “The 202-303 ortho rotational line (5083.93 cm^{−1}, rest) of the 111-100 vibrational hot-band was detected in emission on Sept. 5 and 6 UT at the correct doppler-shifted frequency (5083.81 cm^{−1}) and with the expected intensity. A fluorescence model based on solar pumping in the 000-111 vibrational band, and assuming a rotational temperature of 50 K, leads to a water-production rate of 8 × 10^{27} molecules/s on Sept. 5.5.”

1995 September 18 Daniel W. E. Green